# Post-measurement Quantum Monte Carlo

### v1

@@ III.3 Numerical implementation
For a fixed set of outcomes $\bm{s}$, the sampling algorithm starts from a basis state |$\ket{\alpha}$ as well as the three parts of the operator string $\mathcal{H}_{\bm{a}}$, $\mathcal{P}^{\bm{s}}_{\bm{b}}$, and $\mathcal{P}^{\bm{s}}_{\bm{c}}$, such that the overall configuration is allowed $\braket{\alpha}{\mathcal{P}^{\bm{s}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{s}}_{\bm{c}}}{\alpha}=1$. The basic idea is to propose changes in the operator string and the basis state which take us between different allowed configurations, accepting changes according to their statistical weights (as in any standard Metropolis-Hastings algorithm).
@@
For $B$ consisting only of non-overlapping NN sites, and for outcomes $\bm{s}=0$, the sign of the weight is the parity of the number of type $2$ operators appearing in $\mathcal{P}^{\bm{s}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{s}}_{\bm{c}}$. The condition $\braket{\alpha}{\mathcal{P}^{\bm{s}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{s}}_{\bm{c}}}{\alpha}=1$ implies that this number is even.
@@
For $B$ a set of non-overlapping NNN sites, and for outcomes $\bm{s}=1$, the sign of the weight is the parity of the number of type $2$ operators appearing in $\mathcal{H}_{\bm{a}}$ alone. Spin flips generated by $O^{H}_{(jk),2}$ in $\mathcal{H}_{\bm{a}}$ are along NN $(jk)$ and are associated with negative signs, while spin flips generated by $O^{1}_{(jk)}$ operators in $\mathcal{P}^{\bm{1}}_{\bm{b}}$ and $\mathcal{P}^{\bm{1}}_{\bm{c}}$ are along NNN $(jk)$ and come with positive signs. The condition $\braket{\alpha}{\mathcal{P}^{\bm{s}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{s}}_{\bm{c}}}{\alpha}=1$ then implies that the number of type $2$ operators in $\mathcal{H}_{\bm{a}}$ is even.
This discussion also makes clear that any combination of singlets on NNs and triplets on NNNs is also sign-problem-free, and so can be studied using the post-measurement SSE formulated in the basis |$\ket{\alpha}$.
We now describe our algorithms for sampling operator strings. For ‘singlet measurements’ ($\bm{s}=\bm{0}$ with $B$ a set of NNs) we initialize the $\mathcal{P}^{\bm{0}}_{\bm{b}}$ and $\mathcal{P}^{\bm{0}}_{\bm{c}}$ operator strings with $\bm{b}=\bm{c}=\bm{1}$, i.e. type $1$ operators on all NNs. Recall that, in this setup, we never encounter type 00 operators in the singlet projectors. Therefore, we don’t need to perform ‘diagonal’ moves which exchange type 00 and type $1$ operators in the measurement strings. Instead, in the measurement strings, we only need to focus on off-diagonal updates (which exchange type $1$ and type $2$ operators). Since here the diagonal (type $1$) and off-diagonal (type $2$) operators have weights of the same magnitude, it becomes possible to use highly efficient loop updates [55] of the entire $\mathcal{P}^{\bm{0}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{0}}_{\bm{c}}$ operator string. Such updates satisfy detailed balance and are designed to always give allowed operator strings.
@@
For ‘triplet measurements’ ($\bm{s}=\bm{1}$ with $B$ a set of NNNs) we start with $\mathcal{P}_{\bm{b}}$ and $\mathcal{P}_{\bm{c}}$ consisting of type 00 operators (i.e. identity operators) and then perform both diagonal ($d=0\leftrightarrow d=1$) and off-diagonal ($d=1\leftrightarrow d=2$) updates in these operator strings. Here, when implementing the deterministic loop update, it is important to account for the fact that the diagonal operators $O^{H}_{1,(jk)}$ and $O^{1}_{1,(jk)}$, appearing in $\mathcal{H}_{\bm{a}}$ and the measurement strings, respectively, annihilate different spin configurations: $O^{H}_{1,(jk)}\ket{\uparrow_{j}\uparrow_{k}}=O^{H}_{1,(jk)}\ket{\downarrow_{j}\downarrow_{k}}=0$ while $O^{1}_{1,(jk)}\ket{\uparrow_{j}\downarrow_{k}}=O^{1}_{1,(jk)}\ket{\downarrow_{j}\uparrow_{k}}=0$. The different spin and operator configurations allowed in the SSE are illustrated in Fig. 4.
As a numerical demonstration, we now study the effects of projective singlet and triplet measurements in small systems where we can benchmark our results against exact diagonalization (ED). The two measurement configurations are illustrated in Fig. 5. To test our implementation of $\bm{s}=\bm{0}$ measurements we force $s_{(jk)}=0$ on a non-overlapping set of NNs which cover all but two separated spins. Since at large $\beta$ the initial state $\rho\sim e^{-\beta H}$ approaches the ground state, which is a total singlet, these measurements force the unmeasured spins into a non-local singlet. Therefore $\braket{S^{z}_{j}S^{z}_{k}}^{\bm{0}}=\braket{\bm{S}_{j}\cdot\bm{S}_{k}}/3\to-1/4$ as $\beta\to\infty$. To test our implementation of $\bm{s}=\bm{1}$ measurements we force $s_{(jk)}$ on a non-overlapping set of NNNs which cover every spin; this generates a set of $N/2$ spin-$1$ degrees of freedom whose correlations are inherited from the initial state. The results of our numerical calculations are shown in Fig. 6, and in each case we find quantitative agreement between SSE and ED. In the next section we will study both weak and projective measurements; our results there (e.g. in Fig. 8) demonstrate the accuracy of our weak measurement algorithm.
@@ IV.1 Ordered phase
When we measure the system, the relative contributions of different energy eigenstates to the density matrix are modified, and it is important to understand which eigenstates dominate post-measurement correlations. To address this, we will evaluate $\braket{S^{2}_{\text{total}}}^{\bm{s}}$ in post-measurement states $\rho^{\bm{s}}$. Recall that all eigenstates |$\ket{n}$ of the Hamiltonian $H$ are also eigenstates of the total spin, and that any eigenstate with spin $S_{n}$, i.e. with $S^{2}_{\text{total}}\ket{n}=S_{n}(S_{n}+1)\ket{n}$, belongs to a $(2S_{n}+1)$-fold degenerate multiplet. Moreover, since the measurement operators $P^{\bm{s}}$ are $\mathrm{SU}(2)$ symmetric, they do not mix sectors having different $S^{z}_{\text{total}}$ or $S^{2}_{\text{total}}$ quantum numbers. Denoting by $p^{\bm{s}}_{n}=\braket{n}{P^{\bm{s}}}{n}$ the probability to observe outcomes $\bm{s}$ given eigenstate |$\ket{n}$, the post-measurement expectation value of the total spin $\braket{S^{2}_{\text{total}}}^{\bm{s}}=\text{Tr}[P^{\bm{s}}e^{-\beta H}P^{\bm{s}}S^{2}_{\text{total}}]/\text{Tr}[P^{\bm{s}}e^{-\beta H}]$ is given by
@@
where $H\ket{n}=E_{n}\ket{n}$. Here we have used $\braket{n}{P^{\bm{s}}S^{2}_{\text{total}}P^{\bm{s}}}{n}=S_{n}(S_{n}+1)p^{\bm{s}}_{n}$, which follows from $[P^{\bm{s}},S^{2}_{\text{total}}]=0$. At large inverse temperature $\beta$ we then find
@@
Figure 8: Correlations between unmeasured spins following weak and projective measurements with $\bm{s}=\bm{0}$ on a set of NN pairs of sites excluding 00 and $j$. As the strength of weak measurement $\mu$ is increased, the post-measurement correlations (crosses) converge to the result for projective measurements (solid horizontal lines); see inset. Here $N_{x}=N_{y}$ (legend), $\beta=3\times N_{x}$, and $J_{1}=J_{2}=1$.
This behavior is clear in Fig. 8, where we also test our weak measurement algorithm. We calculate $\braket{\bm{S}_{j}\cdot\bm{S}_{k}}^{\bm{0}}$ for various measurement strengths $\mu$, where the expectation value is with respect to $\rho^{\bm{0}}_{\mu}$. The spatial arrangement of singlet measurements is as depicted in the left panel of Fig. 5, where the two unmeasured spins are maximally separated along the horizontal direction. Our results show that, on increasing the measurement strength $\mu$, the correlation function $\braket{\bm{S}_{j}\cdot\bm{S}_{k}}^{\bm{0}}$ between the unmeasured spins decreases from its ground state value to near $-3/4$, the value for a long-range singlet. However, as the system size is increased with $\beta=3N_{x}$, the post-measurement correlations are indeed weakened.
@@ IV.2 Dimerized phase
Here we show that measurements in the dimerized phase generate SPT order. We consider the Heisenberg model with $J_{2}\gg J_{1}$; in the limit $1/g=J_{1}/J_{2}\to 0$ the ground state becomes a tensor product of singlets on a nonoverlapping set of bonds, and this dimerized phase extends down to $1/g\approx 0.52$. The ground states with $0\leq 1/g\lesssim 1/g_{c}$ are therefore related to the tensor-product ground state at $g=0$ via a finite depth local unitary evolution [57]. Let us denote the pure ground states of this model by |$\ket{\psi_{1/g}}$. Note that since this phase is gapped we have $e^{-\beta H}\sim\ket{\psi_{1/g}}\bra{\psi_{1/g}}$ | already for $\beta$ asymptotically larger than $\log N$.
@@
Each measurement has two possible outcomes. On projecting the measured bond $(x,x+1)$ (with $x$ odd) into a singlet $s_{x,x+1}=0$, we also generate a singlet between spins at sites $x-1$ and $x+2$, i.e. local measurements create a nonlocal singlet. On projecting the measured bond $(x,x+1)$ into a triplet, the two spin-$1/2$ degrees of freedom there are instead merged into a single spin-$1$ degree of freedom. We indicate the effects of these measurements on the right in Fig. 9.
@@
$P^{\bm{s}}\ket{\psi_{0}}\propto\bigotimes_{\sigma}\ket{\text{AKLT}^{\bm{s}}}_{\sigma}$ (36)
where the index $\sigma$ labels the different snakes, and |$\ket{\text{AKLT}^{\bm{s}}}_{\sigma}$ is an AKLT state on the snake $\sigma$, having the locations of its spin-1 degrees of freedom sites fixed by the measurement outcomes $\bm{s}$.
On tuning away from the $01/g=$ 0 limit, the fate of the SPT order is unclear; this is also the case in one dimension, and so in the following we keep our discussion general. We can write the ground state for $1/g<1/g_{c}$ as |$\ket{\psi_{1/g}}=U_{1/g}\ket{\psi_{0}}$, where $U_{1/g}$ is a quasi-local unitary which describes an adiabatic change in the Hamiltonian. The post-measurement state is then $P^{\bm{s}}\ket{\psi_{1/g}}\propto P^{\bm{s}}U_{g}\ket{\psi_{0}}$. It is natural to ask how the string order
@@
behaves for $1/g\neq 0$; here $L^{z}_{a}=P^{\bm{1}}_{(jk)}(S^{z}_{j}+S^{z}_{k})P^{\bm{1}}_{(jk)}$ is a spin-$1$ operator acting on a pair of sites which we here label by $a=(jk)$. For the geometry illustrated in Fig. 10, the index $a$ labels NNNs, all of which feature spin-$1$ degrees of freedom. In Fig. 11 we calculate $O^{\bm{s}}(0,b)$ for the post-measurement state corresponding to $\bm{s}=\bm{1}$ on a nonoverlapping set of NNNs. For $01/g=$ 0 we recover perfect string order as required, and for nonzero $1/g$ within the dimerized phase we find that $O^{\bm{1}}(0,b)$ decays with separation. Here separations are measured along the length of the AKLT chain that would have been created at $01/g=$ 0.
@@ IV.3 Mixed-state entanglement
$\rho^{\bm{s}}_{jk}=\alpha\ket{0,0}\bra{0,0}+\frac{1}{3}(1-\alpha)\sum_{m}\ket{1,m}\bra{1,m},$ (38)
with a parameter $\alpha$ that can be inferred from our numerical calculations via the relation $\alpha=\frac{1}{4}-\braket{\bm{S}_{j}\cdot\bm{S}_{k}}^{\bm{s}}$. In an abuse of notation here we use the notation |$\ket{S,m}$ for the two-qubit states that are eigenstates of $(\bm{S}_{j}+\bm{S}_{k})^{2}$ and $S^{z}_{j}+S^{z}_{k}$, where $j$ and $k$ are in general a pair of separated spins. From Eq. (38) it can be verified that the von Neumann entanglement entropy of either of the spins $j$ and $k$ is maximal regardless of the value of $\alpha$, i.e. $\rho^{\bm{s}}_{j}=\text{Tr}_{k}\rho^{\bm{s}}_{jk}=\frac{1}{2}\mathds{1}$, where $\text{Tr}_{k}$ denotes a partial trace over spin $k$. For $\alpha=1$, as arises for $\beta\to\infty$ in the $\bm{s}=\bm{0}$ measurement scheme of Fig. 6, the spins $j$ and $k$ are maximally entangled with one another. On the other hand, for $\beta=0$, the spins are maximally entangled with the environment, and there is no useful entanglement in the system.
@@
$\mathcal{N}_{jk}=\frac{1}{2}\Big{(}||\rho^{T_{k}}_{jk}||_{1}-1\Big{)}.$ (39)
The negativity vanishes for a separable (i.e. unentangled) mixed states, and therefore can only be nonzero when $j$ and $k$ are entangled. Note, however, that the converse is not true: $j$ and $k$ can be entangled even when $\mathcal{N}_{jk}=0$. For states of the form in Eq. (38) it can be verified that $\mathcal{N}_{jk}\neq 0$ provided $\braket{\bm{S}_{j}\cdot\bm{S}_{k}}<-\frac{3}{20}$.
@@ V.1 Quantum-classical mapping
[…]
@@
[…]
and the sum is over the set of measured pairs of sites $M$. The coordinate $\tau$ should be understood as running around a circle of circumference $\beta+\mu$, with $\tau=-\mu/2$ equivalent to $\tau=\beta+\mu/2$. The Hamiltonian contribution $S_{1}[\Omega]$ only involves an integral over the interval $[\mu/2,\mu/2+\beta]$, while the geometric Berry phase term $S_{2}[\Omega]$ involves an integral over all $\tau\in[0,\beta+\mu]$.
@@ V.3 Ensemble of outcomes
VI Discussion

### v2

@@ III.3 Numerical implementation
[…]
@@
For $B$ consisting only of non-overlapping NN sites, and for outcomes $\bm{s}=0$, the sign of the weight is the parity of the number of type $2$ operators appearing in $\mathcal{P}^{\bm{s}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{s}}_{\bm{c}}$. The condition $\braket{\alpha|\mathcal{P}^{\bm{s}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{s}}_{\bm{c}}|\alpha}=1$ implies that this number is even.
@@
For $B$ a set of non-overlapping NNN sites, and for outcomes $\bm{s}=1$, the sign of the weight is the parity of the number of type $2$ operators appearing in $\mathcal{H}_{\bm{a}}$ alone. Spin flips generated by $O^{H}_{(jk),2}$ in $\mathcal{H}_{\bm{a}}$ are along NN $(jk)$ and are associated with negative signs, while spin flips generated by $O^{1}_{(jk)}$ operators in $\mathcal{P}^{\bm{1}}_{\bm{b}}$ and $\mathcal{P}^{\bm{1}}_{\bm{c}}$ are along NNN $(jk)$ and come with positive signs. The condition $\braket{\alpha|\mathcal{P}^{\bm{s}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{s}}_{\bm{c}}|\alpha}=1$ then implies that the number of type $2$ operators in $\mathcal{H}_{\bm{a}}$ is even.
[…]
We now describe our algorithms for sampling operator strings. For ‘singlet measurements’ ($\bm{s}=\bm{0}$ with $B$ a set of NNs) we initialize the $\mathcal{P}^{\bm{0}}_{\bm{b}}$ and $\mathcal{P}^{\bm{0}}_{\bm{c}}$ operator strings with $\bm{b}=\bm{c}=\bm{1}$, i.e. type $1$ operators on all NNs. Recall that, in this setup, we never encounter type 0 operators in the singlet projectors. Therefore, we don’t need to perform ‘diagonal’ moves which exchange type 0 and type $1$ operators in the measurement strings. Instead, in the measurement strings, we only need to focus on off-diagonal updates (which exchange type $1$ and type $2$ operators). Since here the diagonal (type $1$) and off-diagonal (type $2$) operators have weights of the same magnitude, it becomes possible to use highly efficient loop updates [55] of the entire $\mathcal{P}^{\bm{0}}_{\bm{b}}\mathcal{H}_{\bm{a}}\mathcal{P}^{\bm{0}}_{\bm{c}}$ operator string. Such updates satisfy detailed balance and are designed to always give allowed operator strings.
@@
[…]
As a numerical demonstration, we now study the effects of projective singlet and triplet measurements in small systems where we can benchmark our results against exact diagonalization (ED). The two measurement configurations are illustrated in Fig. 5. To test our implementation of $\bm{s}=\bm{0}$ measurements we force $s_{(jk)}=0$ on a non-overlapping set of NNs which cover all but two separated spins. Since at large $\beta$ the initial state $\rho\sim e^{-\beta H}$ approaches the ground state, which is a total singlet, these measurements force the unmeasured spins into a non-local singlet. Therefore $\braket{S^{z}_{j}S^{z}_{k}}^{\bm{0}}=\braket{\bm{S}_{j}\cdot\bm{S}_{k}}/3\to-1/4$ as $\beta\to\infty$. To test our implementation of $\bm{s}=\bm{1}$ measurements we force $s_{(jk)}$ on a non-overlapping set of NNNs which cover every spin; this generates a set of $N/2$ spin-11 degrees of freedom whose correlations are inherited from the initial state. The results of our numerical calculations are shown in Fig. 6, and in each case we find quantitative agreement between SSE and ED. In the next section we will study both weak and projective measurements; our results there (e.g. in Fig. 8) demonstrate the accuracy of our weak measurement algorithm.
@@ IV.1 Ordered phase
[…]
@@
[…]
@@
Figure 8: Correlations between unmeasured spins following weak and projective measurements with $\bm{s}=\bm{0}$ on a set of NN pairs of sites excluding 0 and $j$. As the strength of weak measurement $\mu$ is increased, the post-measurement correlations (crosses) converge to the result for projective measurements (solid horizontal lines); see inset. Here $N_{x}=N_{y}$ (legend), $\beta=3\times N_{x}$, and $J_{1}=J_{2}=1$.
This behavior is clear in Fig. 8, where we also test our weak measurement algorithm. We calculate $\braket{\bm{S}_{j}\cdot\bm{S}_{k}}^{\bm{0}}$ for various measurement strengths $\mu$, where the expectation value is with respect to $\rho^{\bm{0}}_{\mu}$. The spatial arrangement of singlet measurements is as depicted in the left panel of Fig. 5, where the two unmeasured spins are maximally separated along the horizontal direction. Our results show that, on increasing the measurement strength $\mu$, the correlation function $\braket{\bm{S}_{j}\cdot\bm{S}_{k}}^{\bm{0}}$ between the unmeasured spins decreases from its ground state value to near 3/4-3/4, the value for a long-range singlet. However, as the system size is increased with $\beta=3N_{x}$, the post-measurement correlations are indeed weakened.
@@ IV.2 Dimerized phase
[…]
@@
Each measurement has two possible outcomes. On projecting the measured bond $(x,x+1)$ (with $x$ odd) into a singlet $s_{x,x+1}=0$, we also generate a singlet between spins at sites x1x-1 and $x+2$, i.e. local measurements create a nonlocal singlet. On projecting the measured bond $(x,x+1)$ into a triplet, the two spin-$1/2$ degrees of freedom there are instead merged into a single spin-$1$ degree of freedom. We indicate the effects of these measurements on the right in Fig. 9.
@@
[…]
@@
behaves for $1/g\neq 0$; here $L^{z}_{a}=P^{\bm{1}}_{(jk)}(S^{z}_{j}+S^{z}_{k})P^{\bm{1}}_{(jk)}$ is a spin-$1$ operator acting on a pair of sites which we here label by $a=(jk)$. For the geometry illustrated in Fig. 10, the index $a$ labels NNNs, all of which feature spin-$1$ degrees of freedom. In Fig. 11 we calculate $O^{\bm{s}}(0,b)$ for the post-measurement state corresponding to $\bm{s}=\bm{1}$ on a nonoverlapping set of NNNs. For $1/g=0$ we recover perfect string order as required, and for nonzero $1/g$ within the dimerized phase we find that $O^{\bm{1}}(0,b)$ decays with separation. Here separations are measured along the length of the AKLT chain that would have been created at $1/g=0$.
@@ IV.3 Mixed-state entanglement
[…]
@@
[…]
The negativity vanishes for separable (i.e. unentangled) mixed states, and therefore can only be nonzero when $j$ and $k$ are entangled. Note, however, that the converse is not true: $j$ and $k$ can be entangled even when $\mathcal{N}_{jk}=0$. For states of the form in Eq. (38) it can be verified that $\mathcal{N}_{jk}\neq 0$ provided $\braket{\bm{S}_{j}\cdot\bm{S}_{k}}<-\frac{3}{20}$.
@@ V.1 Quantum-classical mapping
$S_{1}[\Omega]=\sum_{(jk)\in\text{NN}}J_{(jk)}\int_{0}^{\beta}d\tau\,\bm{\Omega}_{j}(\tau)\cdot\bm{\Omega}_{k}(\tau),$ (42)
@@
$S[\Omega]=S_{1}[\Omega]+S_{2}[\Omega]-\int_{-\mu/2}^{\mu/2}d\tau\sum_{(jk)\in M}\bm{\Omega}_{j}\cdot\bm{\Omega}_{k},$ (45)
and the sum is over the set of measured pairs of sites $M$. The coordinate $\tau$ should be understood as running around a circle of circumference $\beta+\mu$, with $\tau=-\mu/2$ equivalent to $\tau=\beta+\mu/2$. The Hamiltonian contribution $S_{1}[\Omega]$ only involves an integral over the interval $\mu/2,\mu/2+\beta]$, while the geometric Berry phase term $S_{2}[\Omega]$ involves an integral over all $\tau\in[0,\beta+\mu]$.
@@ V.3 Ensemble of outcomes
VI Single-body measurements
In this section we study the effects of measuring local spin operators. Because the unmeasured SSE for the Heisenberg model has no sign problem when working in a fixed site-local basis, measurements in this basis do not introduce a sign problem in the post-measurement SSE. It is important to note that this feature is not specific to the Heisenberg model: If there is a basis in which the unmeasured SSE is sign-free, the effects of measurements in that basis can be studied using a sign-free post-measurement SSE, regardless of the measurement outcomes. Conditional correlations in all post-measurement states can then be evaluated efficiently using our algorithm.
In the present setting, measurements of local spins break the $\mathrm{SU}(2)$ symmetry of the density matrix and have qualitatively different effects to the symmetric measurements studied in previous sections. Let us first divide the spins into two sets $A$ and $B$: we will projectively measure $S^{z}_{j}$ for spins $j$ in set $B$, and ask how these measurements affect correlations between spins in set $A$. In order to do this we need to (i) sample outcomes of measurements, (ii) calculate post-measurement correlations conditioned on specific sets of measurement outcomes, and (iii) construct a suitable average over these post-measurement correlations.
[…]
Each observed set of measurement outcomes $\alpha_{j}$ defines a set of projection operators $P^{\alpha_{j}}=\frac{1}{2}(1+4\alpha_{j}S^{z}_{j})$, such that the overall projector is
$P^{\alpha_{B}}=\bigotimes_{j\in B}P^{\alpha_{j}}.$ (49)
The partition function which can be used to generate post-measurement correlations is then
$Z^{\alpha_{B}}$ $=\text{Tr}[P^{\alpha_{B}}e^{-\beta H}P^{\alpha_{B}}]$ (50)
$=\sum_{\alpha_{A}}\braket{\alpha_{A}\alpha_{B}|e^{-\beta H}|\alpha_{A}\alpha_{B}},$
and we will only compute correlations between sites $j,k\in A$. Expanding $e^{-\beta H}$ in powers of $\beta$, we can define a post-measurement SSE which samples operator strings as well as configurations $\alpha_{A}$ of the $A$ sites that are consistent with the fixed configuration $\alpha_{B}$ of the $B$ sites. To implement this algorithm, during the loop update step we construct closed loops around the $B$ sites at $\tau=0$, thereby ensuring that their states are not updated; we illustrate these loops in Fig. 15. Ergodicity is preserved over the unmeasured $A$ spins, allowing us to efficiently sample the ensemble of measured states.
Figure 15: Freezing loop update for single-body $S^{z}_{j}$ projectors $\frac{1}{2}(1\pm 2S^{z}_{j})$ in the Heisenberg model. A closed loop is initiated at each projector using the deterministic loop [27] update scheme. All operator legs that belong to the constructed loop are frozen—i.e., excluded from further updates—during the off-diagonal update. Frozen legs are indicated in maroon and the arrows show the direction of the loop. Projectors are shown as open red rectangles, and here even sites belong to the set $B$.
Using these methods, we can calculate post-measurement correlations in states $\rho^{\alpha_{B}}\propto P^{\alpha_{B}}e^{-\beta H}P^{\alpha_{B}}$, e.g. $\braket{S^{z}_{j}S^{z}_{k}}^{\alpha_{B}}\equiv\text{Tr}[\rho^{\alpha_{B}}S^{z}_{j}S^{z}_{k}]$. We now seek a suitable average over correlation functions that can detect the effects of measurements. Denoting the probability to observe $\alpha_{B}$ by $p^{\alpha_{B}}$, recall that
$\sum_{\alpha_{B}}p^{\alpha_{B}}\braket{S^{z}_{j}S^{z}_{k}}^{\alpha_{B}}=\braket{S^{z}_{j}S^{z}_{k}},$ (51)
i.e. averaging post-measurement correlations $\braket{S^{z}_{j}S^{z}_{k}}^{\alpha_{B}}$ according to the Born rule ‘washes out’ the nonlocal effects of our measurements. A simple probe of the effects of the measurements is the measurement-averaged connected correlation function
$C_{jk}=\sum_{\alpha_{B}}p^{\alpha_{B}}\Big{[}\braket{S^{z}_{j}S^{z}_{k}}^{\alpha_{B}}-\braket{S^{z}_{j}}^{\alpha_{B}}\braket{S^{z}_{k}}^{\alpha_{B}}\Big{]}.$ (52)
The second term in the brackets is nonlinear in the post-measurement density matrix, and is generically nonzero. Note that without measurements $\braket{S^{z}_{j}}=0$, so this connected correlation function simplifies to $\braket{S^{z}_{j}S^{z}_{k}}$.
In practice, we calculate $C_{jk}$ as follows. Initially, we run an unmeasured SSE simulation (Sec. II.3) and sample configurations $\alpha_{B}$ of the $B$ spins at large intervals along the Markov chain; we denote the number of different samples used by $N_{B}$. Then for each of the $N_{B}$ samples, we require estimates for both $\braket{S^{z}_{j}S^{z}_{k}}^{\alpha_{B}}$ and $\braket{S^{z}_{j}}^{\alpha_{B}}\braket{S^{z}_{k}}^{\alpha_{B}}$. To evaluate the latter, we run two statistically independent post-measurement SSEs: If the errors in our estimates for $\braket{S^{z}_{j}}^{\alpha_{B}}$ and $\braket{S^{z}_{k}}^{\alpha_{B}}$ are of order $\epsilon$, the error in our estimate for $\sum_{\alpha_{B}}p^{\alpha_{B}}\braket{S^{z}_{j}}^{\alpha_{B}}\braket{S^{z}_{k}}^{\alpha_{B}}$ is then of order $\epsilon/\sqrt{N_{B}}$. On the other hand, if our estimates for $\braket{S^{z}_{j}}^{\alpha_{B}}$ and $\braket{S^{z}_{k}}^{\alpha_{B}}$ were correlated, there would be a drift in our estimate for $\sum_{\alpha_{B}}p^{\alpha_{B}}\braket{S^{z}_{j}}^{\alpha_{B}}\braket{S^{z}_{k}}^{\alpha_{B}}$ scaling as $\epsilon^{2}/N_{B}$; by using two independent Markov chains, we avoid this drift. From the post-measurement SSEs we also obtain estimates for $\braket{S^{z}_{j}S^{z}_{k}}^{\alpha_{B}}$, and we average these estimates over observed $\alpha_{B}$ to estimate $\sum_{\alpha_{B}}p^{\alpha_{B}}\braket{S^{z}_{j}S^{z}_{k}}^{\alpha_{B}}$. Agreement between the results of this calculation, and the correlations $\braket{S^{z}_{j}S^{z}_{k}}$ in the unmeasured state, serve as a simple consistency check.
We now calculate $C_{jk}$ for the Heisenberg AFM tuned to the quantum critical regime, with $\alpha_{B}$ the entire $B$ sublattice, and with spins $j,k\in A$ separated along the principal diagonal of the lattice. The geometry of the set of measured spins $B$ and the principal diagonal along which we measure $C_{jk}$ is shown in Fig. 16. To understand what we should expect here, recall that the scaling dimension $\Delta$ of the spin operator at an $\mathrm{O}(3)$ critical point is close to its mean field value of $1/2$, i.e. two point correlations in the unmeasured system decay algebraically with an exponent close to unity. Following Ref. [16], for a quantum critical system in $d$ spatial dimensions, a weak measurement of an operator with scaling dimension $\Delta$ constitutes a relevant perturbation if $d>2\Delta$, which is the case here. Since weak measurements of the spin operator are relevant, it is natural to expect that projective measurements destroy long-range correlations. Therefore, we expect that while the algebraic decay of the algebraic decay of $\braket{S^{z}_{j}S^{z}_{k}}$ in the unmeasured system will be replaced by a rapid decay of $C_{jk}$ in the post-measurement state. Our numerical results, shown in Fig. 17, are consistent with this.
In summary, here we have shown that the post-measurement SSE can be used to evaluate correlations in states generated by the measurement of local spin operators in a fixed basis. Here, because there is no sign problem, our algorithm can be used to construct probes of post-measurement correlations [36, 37] relevant to experiments in quantum simulators.
Figure 16: Figure illustrating single site $S^{z}_{j}$-measurement setup on square lattice for AFM Heisenberg model. Spins on the $B$ sublattice (indicated with squares) are measured while spins on the AA sublattice are not measured. The diagonal enclosed within the rectangular box indicates the direction along which we measure the measurement averaged connected correlation function $C_{0k}$ given by Eq. (52)
Figure 17: Connected spin correlations along the principal diagonal of a square lattice in the quantum critical regime with ($\mu=0$) the $B$ sublattice unmeasured and ($\mu=\infty$) the $B$ sublattice projectively measured in $S^{z}_{j}$ basis. Here $N_{x}=N_{y}=16$, $\beta=3N_{x}=48$, $J_{1}=1$, and $J_{2}=1$.
VII Discussion